Circular No. 6851

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980326 G. Celidonio and A. Coletta, BeppoSAX Science Operation Center,

Telespazio, Rome; M. Feroci, L. Piro, and P. Soffitta, Istituto Astrofisica Spaziale, CNR, Rome; J. in ’t Zand, Space Research Organization of the Netherlands (SRON), Utrecht; J. Muller, BeppoSAX Science Data Center, Rome and SRON, and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, Bologna, report: “The BeppoSAX Gamma Ray Burst Monitor was triggered on Mar. 26.888125 UT, coincident with BATSE trigger 6660. A quick-look analysis shows a single-peak γ-ray burst lasting ∼ 5 s and having a peak flux of 500 counts/s in the range 40–700 keV. The γray burst was also detected by unit 1 of the Wide Field Cameras (WFC1) with a peak flux of ∼ 4 Crab in the range 2–26 keV. The position from a quick-look analysis is α = 8^{h}36^{m}26^{s}, δ = −18^{o}53′.0 (equinox 2000.0). The accuracy of the satellite attitude was not optimum, and we adopt an error radius of 8′. Due to sun constraints, the GRB location cannot be presently observed by BeppoSAX Narrow Field Instruments.”

COMET 55P/TEMPEL-TUTTLE P. Lamy, Laboratoire d’Astronomie Spatiale, and his team report the

detection of the nucleus of comet 55P with the Hubble Space Telescope (+ WFPC2) on Jan. 9 at Δ = 0.46 AU: “After removing the faint contribution of the coma to the central pixels, the magnitude of the nucleus was R = 16.62, showing no variation over 10 hr. For a geometric albedo of 0.04 and a phase coefficient of 0.04 mag/deg, the mean effective radius is 1.8 km, the largest uncertainty coming from the phase law. The red color of the nucleus is characterized by a normalized reflectivity gradient of 0.16 percent/nm.”

Visual m_{1} estimates: Mar. 1.75 UT, 8.7 (V. Znojil, Brno, Czech R., 25×100 binoculars); 18.24, 9.7: (N. Biver, Honolulu, HI, 0.26-m reflector).

NOVA SAGITTARII 1998 B. A. Skiff, Lowell Observatory, reports that examination of the orig-

inal Palomar Sky Survey prints show nothing definite at the location of N Sgr 1998, indicating a lower limit on the outburst amplitude of at least 11 (and perhaps 12) magnitudes.

Additional visual magnitude estimates: Mar. 25.12 UT, 8.4 (L. Kiss, Szeged, Hungary); 26.141, 8.1 (K. Hornoch, Lelekovice, Czech Republic); 26.66, 8.2 (F. Farrell, Christies Beach, S. Australia); 27.17, 8.7 (A. Baransky, Kiev, Ukraine); 27.215, 8.4 (A. Pereira, Cabo da Roca, Portugal).

1998 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green